Many years ago, Topnotch Knives issued a zero coupon bond with a $1,000 face value. The...

  • Many years ago, Topnotch Knives issued a zero coupon bond with a $1,000 face value. The bond matures in three years. If the current mar- ket rate on similar bonds is 11 percent, (a) what is the bond’s current value? Suppose the market rate stays at 11 percent for the next three years. What (b) current yield and (c) capital gains yield will bondholders receive each year during the remainder of the bond’s life?
Answer #1

It is given that Zero coupon bond TK bond face value is $1,000. Bond is matured in three years. Market rate of the bond is 1

= (0) +1,000(0.73121) = 731.21 Hence, current value of bond is $731.21

Current yield for year 1 can be calculated as follows: Coupon interest ratex Maturity value Current yield = Previous year bon

= 0 Hence, current yield for year 3 is o percent Capital gain for year 1 can be calculated as follows: (First year year price

811.64-731.21 731.21 = 0.11 Hence, capital gain for year 1 is 11 percent. Total yield for year 1 can be calculated as follows

Capital gain for year 2 can be calculated as follows: (First year year price (*(1 + Yield to maturity) Time period (First yea

Hence, capital gain for year 2 is 11 percent Total yield for year 2 can be calculated as follows: Total yield = Current yield

Capital gain for year 3 can be calculated as follows: (First year year price (*(1 + Yield to maturity) Time period First year

Hence, capital gain for year 3 is 11 percent Total yield for year 3 can be calculated as follows: Total yield = Current yield